OpenStudy (anonymous):

how do you prove that a line parallel to one side of a triangle divides the other two sides proportionally? And...

OpenStudy (anonymous):

how do you prove the base angles are congruent in an isosceles trapezoid.

OpenStudy (anonymous):

And What is the difference between the Pythagorean Theorem and the Converse of the Pythagorean Theorem??

OpenStudy (anonymous):

A kite is a quadrilateral that has diagonals that are perpendicular bisectors of one another. how would I prove that the adjacent sides are congruent

OpenStudy (anonymous):

Please help..

OpenStudy (anonymous):

@zepdrix @texaschic101

OpenStudy (anonymous):

@perl @YungKhris @StudyGurl14 @SithsAndGiggles @linn99123

OpenStudy (anonymous):

@Destinymasha @prowrestler @KendrickLamar2014

OpenStudy (studygurl14):

What is the difference between the Pythagorean Theorem and the Converse of the Pythagorean Theorem Phytahorean Theorem: If a triangle is a right triangle, sides a, b, and c, with c being the hypotenuse, then... \[a^2+b^2=c^2\] Converse: If a triangles' sides a, b, and c, satisfy the equation a^2+b^2=c^2, the triangle is a rigth traingle

OpenStudy (kendricklamar2014):

Im not that good at Math, sorry :(